# Predicting perturbation targets with causal differential networks

### v1

@@ Abstract
Rationally identifying variables responsible for changes to a biological system can enable myriad applications in disease understanding and cell engineering. From a causality perspective, we are given two datasets generated by the same causal model, one observational (control) and one interventional (perturbed). The goal is to isolate the subset of measured variables (e.g. genes) that were the targets of the intervention, i.e. those whose conditional independencies have changed. Knowing the causal graph would limit the search space, allowing us to efficiently pinpoint these variables. However, current algorithms that infer causal graphs in the presence of unknown intervention targets scale poorly to the hundreds or thousands of variables in biological data, as they must jointly search the combinatorial spaces of graphs and consistent intervention targets. In this work, we propose a causality-inspired approach for predicting perturbation targets that decouples the two search steps. First, we use an amortized causal discovery model to separately infer causal graphs from the observational and interventional datasets. Then, we learn to map these paired graphs to the sets of variables that were intervened upon, in a supervised learning framework. This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets, each with thousands of measured variables. We also demonstrate significant improvements over six causal discovery algorithms in predicting intervention targets across a variety of tractable, synthetic datasets.
@@ 1 Introduction
Cells form the basis of biological systems, and they take on a multitude of dynamical states throughout their lifetime. In addition to natural factors like cell cycle, external perturbations (e.g. drugs, gene knockdown) can alter a cell’s state. While perturbations can affect numerous downstream variables, identifying the root causes, or targets, that drive these transitions has vast therapeutic implications, from cellular reprogramming (Cherry & Daley, 2012) to mechanism of action elucidation (Schenone et al., 2013). Large-scale experiments (Replogle et al., 2022) have attempted to map the effects of perturbing individual genes on single cells, and in-silico approaches (Roohani et al., 2023; Lotfollahi et al., 2023) have been designed for the “forward” inference task of predicting the post-perturbation expression of each gene. In principle, these models can be used within an active learning framework to suggest perturbation targets (Huang et al., 2023; Zhang et al., 2023a). However, the number of inference calls required scales exponentially with the size of the candidate set (e.g. drugs with off-target effects), rendering these approaches impractical for larger search spaces. These models’ predictive performance as oracles has also been called into question by subsequent works (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024), highlighting the difficulty of this forward approach. Finally, there are limited training data for combinatorial perturbations: the most widely used dataset only contains around a hundred pairs (Norman et al., 2019).
The “reverse” strategy – of directly predicting targets from perturbation data – alleviates these problems to an extent. Towards this end, it is common to assume (sometimes implicitly) that the data were generated by a structured causal model, and the perturbation targets can be directly inferred from the changes to this model. While the effects of perturbations are highly contextual, specific to the cell lines in which they are applied (Nadig et al., 2024), existing methods rely on non-specific, data-mined knowledge graphs as the backbones for their causal structures (Cosgrove et al., 2008; Gonzalez et al., 2024). It is hard to quantify the extent to which these graphs are relevant to each setting, as the quality and focus of assays have changed over time, and interactions present in one cell type may not exist in another (Huttlin et al., 2021).
Our goal is to predict the perturbation targets that translate one set of cells into another, while inferring the underlying causal structures directly from the data. We formalize this task within a causality framework. Specifically, given an observational dataset (e.g. gene expression of control cells) and an interventional dataset (e.g. perturbed cells) generated by the same system (e.g. cell type), we would like to identify the variables (e.g. genes) whose data-generating mechanisms are different between the two settings. Causal discovery algorithms have been designed to solve this exact problem, on interventional data with unknown targets (Squires et al., 2020; Brouillard et al., 2020; Hägele et al., 2023). These algorithms operate jointly over the spaces of graphs and intervention target sets. Due to this large search space, these algorithms are unable to scale to the thousands of variables and hundreds of thousands of observations in modern transcriptomics data, while simultaneously being robust to as few as tens of observations per interventional setting (Nadig et al., 2024).
To address unknown intervention target prediction in context of biological perturbations, we propose causal differential networks (Cdn). Drawing upon recent advances in scalable and data-efficient causal discovery (Wu et al., 2024), Cdn first “featurizes” the observational and interventional datasets in terms of their predicted causal graphs, using a pretrained causal discovery module. These paired, edge-level representations are then used as input to an axial attention-based classifier (Ho et al., 2020), which is supervised by ground truth intervention targets (on either real or synthetic data). We thoroughly evaluate Cdn on both real transcriptomic data and in synthetic settings. Cdn outperforms the state-of-the-art in perturbation modeling on the five largest Perturb-seq datasets at the time of publication (CRISPRi,  Replogle et al. (2022); Nadig et al. (2024)), and two Sci-Plex datasets (chemical perturbations, McFaline-Figueroa et al. (2024)), without using any external knowledge. In fact, Cdn is the only model that consistently ranks the ground truth perturbation targets higher than would be expected by random. Moreoever, Cdn is able to maintain decent performance even on unseen cell lines, demonstrating its potential to generalize. Finally, on synthetic settings, Cdn outperforms causal discovery algorithms for unknown intervention targets. To conclude, our contributions are three-fold.
  1. 1.
    We propose causal differential networks (Cdn), a causal discovery-based approach for perturbation target prediction on transcriptomic data, which achieves the state-of-the-art on genetic and chemical perturbation datasets.
  2. 2.
    We demonstrate that Cdn outperforms current causal discovery approaches for unknown intervention target prediction for both hard and soft interventions in diverse synthetic data.
  3. 3.
    Finally, we have prepared seven high-quality transcriptomics datasets, to serve as a benchmark for future machine learning studies.
@@
Figure 1: (A) Biological applications. (B) Differences between observation dataset $D_{\text{obs}}$ and interventional dataset $D_{\text{int}}$ can be attributed to changes in the underlying causal model $G$. (C) We use an amortized causal discovery module to obtain features that represent $G_{\text{obs}}$ and $G_{\text{int}}$. Then we train a supervised classifier to predict the unknown intervention targets.
@@ 2 Background and related work
2.1 Identifying perturbation targets
The rise of large-scale perturbation screens (Replogle et al., 2022) has enabled machine learning approaches for predicting perturbation targets. One line of work focuses on active learning to reduce experimental costs (Zhang et al., 2023a; Huang et al., 2023), but these models require inference calls to an oracle model for each candidate perturbation. Given that human cells express over ten thousand genes, and drugs may impact multiple targets, the search space is enormous. Moreover, subsequent works have found that naive baselines (mean of the training set, linear regression) outperform these approaches at their intended task (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024). In contrast, perturbation targets can be predicted directly from data, using a variety of domain knowledge (Cosgrove et al., 2008; Gonzalez et al., 2024; Roohani et al., 2024). However, these external data were collected from highly inhomogenous sources, which may be inconsistent with the data at hand.
2.2 Causal discovery with unknown intervention targets
Perturbation target prediction can be formalized as a causal discovery problem, using data generated under interventions with unknown targets. Specifically, a causal graphical model is a directed graph $G=(V,E)$, where nodes $i\in V$ map to random variables $X_{i}\in X$, and edges $(i,j)\in E$ represent relationships from $X_{i}$ to $X_{j}$. There are a number of common assumptions that relate $G$ to the data distribution $P_{X}$, which we defer to works such as Spirtes et al. (2001); Yang et al. (2018); Zhang et al. (2023b), since the identifiability of any particular system is not the focus of this paper.
Data generated directly from $P_{X}$ are known as observational data. A causal graphical model allows us to perform interventions by assigning new conditionals
@@ 2.2 Causal discovery with unknown intervention targets
This task has been well-studied in the causality literature, often as a sub-problem or setting while inferring graph $G$ from $D_{\text{obs}}$ and (multiple) $D_{\text{int}}$. Ke et al. (2019) first proposed a discovery algorithm for unknown interventions, on discrete data of up to 50 variables. Jaber et al. (2020) provides identifiability guarantees for soft interventions with unknown targets and proposes a constraint-based algorithm, which requires an exponential number of conditional independence tests with respect to the number of variables. Continuous causal discovery algorithms that are consistent despite unknown interventions include Dcdi (Brouillard et al., 2020) and BaCaDI (Hägele et al., 2023), which treat intervention target identities as learnable parameters of a generative model over $P_{X}$. Dcdi uses the Gumbel-Softmax approximation (Jang et al., 2017) to learn these categorical variables, while Bacadi takes a Bayesian approach, with gradient-based methods for efficient posterior estimation (Liu & Wang, 2016). Finally, Varici et al. (2022) and Yang et al. (2024) frame intervention target prediction as a separate task, similar to this paper. However, the former strictly assumes linearity, while the latter requires data from multiple environments, which are not available here.
2.3 Amortized causal discovery algorithms
The majority of causal discovery algorithms operate on one or more datasets $D$, which correspond to a single graph $G$ (Spirtes et al., 1995; Hauser & Bühlmann, 2012; Mooij et al., 2020; Brouillard et al., 2020). These algorithms must be fit or run from scratch for each data distribution – a potential challenge in low-resource scenarios, where there are too few observations of too many variables. Recently, amortized causal discovery algorithms have addressed causal structure learning as a supervised machine learning problem. Using large numbers of synthetic datasets, generated by (known) synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Ke et al., 2022; Lorch et al., 2022; Wu et al., 2024). Two immediate benefits are fast runtimes and robustness to aleatoric noise. In this work, we leverage the framework proposed by Wu et al. (2024), as it has been benchmarked on large, transcriptomics data (Chevalley et al., 2022).
@@ 3 Methods
Our model is composed of two modules: a frozen causal featurizer, and a learned differential network, which we refer to jointly as Cdn (Figure 2). The causal featurizer is a pretrained, amortized causal discovery model (Wu et al., 2024) that runs efficiently on up to a thousand nodes (Section 3.1), while the differential network is an axial-attention based classifier (Ho et al., 2020) that predicts the nodes whose data-generating mechanisms have changed (Section 3.2).
3.1 Causal featurizer
The causal featurizer takes as input datasets $D_{\text{obs}}$ and $D_{\text{int}}$ containing samples of $N$ variables, generated from observational and interventional settings. For example, $D_{\text{obs}}$ may correspond the gene expression matrix of non-targeting control cells; $D_{\text{int}}$ is the gene expression of the same type of cells, subject to a perturbation with an unknown target; and $N$ is the number of genes detected.
We use an amortized causal discovery model to obtain features $h_{\text{obs}},h_{\text{int}}\in\mathbb{R}^{N\times N\times d}$, which reflect the $N\times N$ adjacency matrices of causal graphs that may have generated each dataset. Specifically, to featurize each dataset, we follow the three steps proposed in Wu et al. (2024).

### v2

# Identifying biological perturbation targets through causal differential networks

@@ Abstract
Identifying variables responsible for changes to a biological system enables applications in drug target discovery and cell engineering. Given a pair of observational and interventional datasets, the goal is to isolate the subset of observed variables that were the targets of the intervention. Directly applying causal discovery algorithms is challenging: the data may contain thousands of variables with as few as tens of samples per intervention, and biological systems do not adhere to classical causality assumptions. We propose a causality-inspired approach to address this practical setting. First, we infer noisy causal graphs from the observational and interventional data. Then, we learn to map the differences between these graphs, along with additional statistical features, to sets of variables that were intervened upon. Both modules are jointly trained in a supervised framework, on simulated and real data that reflect the nature of biological interventions. This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets. We also demonstrate significant improvements over current causal discovery methods for predicting soft and hard intervention targets across a variety of synthetic data.
perturbation, transcriptomics, causality
@@ 1 Introduction
Cells form the basis of biological systems, and they take on a multitude of dynamical states throughout their lifetime. In addition to natural factors like cell cycle, external perturbations (e.g. drugs, gene knockdown) can alter a cell’s state. While perturbations can affect numerous downstream variables, identifying the root causes, or targets, that induce these transitions has vast therapeutic implications, from cellular reprogramming (Cherry & Daley, 2012) to mechanism of action elucidation (Schenone et al., 2013).
Machine learning methods have primarily been designed to infer the effects of perturbations on cells, with the goal of generalizing to unseen perturbations (Roohani et al., 2023), or unseen cell distributions (Lotfollahi et al., 2023; Bunne et al., 2023). In principle, these models can be used within an active learning framework to discover intervention targets (Huang et al., 2023; Zhang et al., 2023a). However, the number of candidate sets scales exponentially with the number of variables under consideration, and there are limited training data for combinatorial perturbations ($<150$ pairs in Norman et al. (2019)). Thus, this option may be impractical for larger search spaces.
Alternatively, some methods predict targets by comparing pairs of “before” and “after” cell states, similar to this paper. They do so by attributing observed differences to sparse, mechanistic changes. Specifically, relationships between variables are represented through a graph, e.g. a gene regulatory network (Babu et al., 2004). Existing methods often use data-mined relations as priors for this graph, and search for modifications to the set of edges that best explain the interventional data (Cosgrove et al., 2008; Gonzalez et al., 2024). However, biological knowledge graphs are incomplete and noisy priors. They contain heterogeneous, potentially conflicting information, since the reported findings were observed under diverse contexts (Ashburner et al., 2000). Moreover, different cell populations may exhibit distinct gene regulatory behavior (Huttlin et al., 2021), leading to incomplete knowledge, especially of less well-characterized systems.
In this paper, we propose causal differential networks (Cdn): a causality-inspired approach to identify variables that drive desired shifts in cell state, while estimating their mechanistic structure directly from data. Given a pair of observational and interventional datasets, we train a causal structure learner to predict causal graphs that could have generated each dataset. The pair of graphs is input to an attention-based classifier, which predicts whether each variable was subject to intervention. To address the challenges of data noise and sparsity, and to simulate the structure of biological interventions, these models are trained jointly in a supervised (amortized) framework (Ke et al., 2022; Petersen et al., 2023), over thousands of synthetic or real datasets.
We evaluate Cdn on real transcriptomic data and synthetic settings. Cdn outperforms the state-of-the-art in perturbation modeling (deep learning and statistical approaches) on the five largest Perturb-seq datasets at the time of publication  (Replogle et al., 2022; Nadig et al., 2024), without using any external knowledge. Furthermore, Cdn generalizes with minimal performance drop to unseen cell lines, which have different supports (genes), causal mechanisms (gene regulatory networks), and data distributions. Finally, on synthetic settings, Cdn outperforms causal discovery approaches for estimating unknown intervention targets.
@@
Figure 1: (A) Biological applications. (B) Differences between observation dataset $D_{\text{obs}}$ and interventional dataset $D_{\text{int}}$ can be attributed to changes in the underlying causal mechanisms. (C) We use an amortized causal discovery model to predict $G_{\text{obs}}$ and $G_{\text{int}}$, whose hidden representations are input to a perturbation target classifier. Both modules are jointly trained.
@@ 2 Background and related work
2.1 Problem statement
We consider a system of $N$ random variables $X_{i}\in X$, whose relationships are represented by directed graph $G=(V,E)$, where nodes $i\in V$ map to variables $X_{i}$, and edges $(i,j)\in E$ to relationships from $X_{i}$ to $X_{j}$. Samples from the data distribution $P_{X}$ are known as observational. We can perform interventions by assigning new conditionals
@@ 2.1 Problem statement
In the context of molecular biology, a perturbation is a hard or soft intervention, and $I$ is the set of perturbation targets, where $\|I\|\geq 1$.
2.2 Modeling perturbations
Perturbation experiments result in datasets of the form $D_{\text{obs}},\{(I^{k},D_{\text{int}}^{k})\}$, where $D_{\text{obs}}\sim P_{X}$, $D_{\text{int}}^{k}\sim\tilde{P}^{k}_{X}$. In this paper, we focus on the transcriptomic data modality, where variables $X_{i}$ represent the levels of gene $i$; the number of variables $N$ ranges from hundreds to thousands; $D_{\text{obs}}$ contains thousands of samples; and each $D_{\text{int}}$ contains tens to hundreds of samples (Replogle et al., 2022).
Each experiment is conducted in a single cell line, which represents a distinct set of variables $X$, distribution $P_{X}$, and graph $G$. These three objects may share similarities across cell lines, as certain genes or pathways are essential for survival. However, the empirical discrepancy is quite large between experiments. For example, Nadig et al. (2024) reports that the median correlation in log-fold change of genes under the same perturbation, across two experiments in the same cell line, by the same lab, was 0.16.
There are two directions for modeling perturbations. To reduce experimental costs, which scale with the number of cell lines and perturbations, one option is to infer the post-intervention distribution of cells. That is, given $D_{\text{obs}}\sim P_{X}$ and $I$, the goal is to infer $\tilde{P}_{X}$. Works such as scGen (Lotfollahi et al., 2019), CPA (Lotfollahi et al., 2023), ChemCPA (Hetzel et al., 2022), and CellOT (Bunne et al., 2023) aim to generalize to new settings $P_{X}$, e.g. unseen cell lines or patients. Alternatively, models like GEARS (Roohani et al., 2023), AttentionPert (Bai et al., 2024), and GIM (Schneider et al., 2024) are designed to generalize to unseen perturbations. These approaches do not explicitly predict targets, but if the space of perturbations $\mathcal{I}$ can be efficiently enumerated, these models are useful for comparing different $I$ (Zhang et al., 2023a; Huang et al., 2023). For example, given an oracle $\phi:D_{\text{obs}},I\to\tilde{P}$, the $I^{*}$ associated with $\tilde{P}^{*}$ is
$I^{*}=\arg\min_{I\in\mathcal{I}}d(\tilde{P}^{*},\phi(D_{\text{obs}},I))$ (3)
for some dissimilarity $d$. However, recent works have reported that the predictive ability of current models does not exceed that of naive baselines (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024), reflecting potential barriers to this approach.
Perturbation datasets can also be used to train models for predicting targets of interest, as is the focus of this work. A common strategy is to relate observed changes (e.g. disease, gene levels) to mechanistic explanations (e.g. pathways), in which the targets are involved. To draw this connection, current approaches use external information such as knowledge graphs (Cosgrove et al., 2008; Gonzalez et al., 2024) and natural language (Roohani et al., 2024). However, these information are collected from highly inhomogenous sources (Ashburner et al., 2000), and it is difficult to quantify their correctness and completeness in each new setting. In contrast to these approaches, we choose to infer the mechanistic structure directly from data.
2.3 Causal discovery
The task of inferring causal mechanisms from data is known as causal discovery, or causal structure learning (Spirtes et al., 2001). In the simplest setting, the goal is to infer $G$ from $D_{\text{obs}}$. The earliest works were discrete search algorithms, which operated over the combinatorial space of potential graphs (Spirtes et al., 1995; Tian & Pearl, 2001; Chickering, 2002; Shimizu et al., 2006). More recently, a number of works have reframed the task as a constrained optimization over continuous adjacency matrices (Zheng et al., 2018; Lachapelle et al., 2020; Sanchez et al., 2023).
If data from multiple environments are available, they can provide additional identifiability regarding the system (Jaber et al., 2020; Zhang et al., 2023b). Differences between these environments, e.g. intervention targets $I$, are either given as input or part of the inference task. As in the observational case, discovery algorithms can be categorized into discrete search (Hauser & Bühlmann, 2012; Ghassami et al., 2018; Ke et al., 2019; Huang et al., 2020; Mooij et al., 2020) or continuous optimization (Brouillard et al., 2020; Hägele et al., 2023). Additionally, Varici et al. (2022) and Yang et al. (2024) frame intervention target prediction as its own task, similar to this paper. However, the former strictly assumes linearity, while the latter requires at least as many environments as $\|I\|$ (we only consider before and after perturbation). Finally, Dci (Belyaeva et al., 2021) infers edge-level differences between linear Gaussian models, rather than identifying the full causal structure. Our work is also motivated by the idea that differences between summary statistics (e.g. regression coefficients in Dci) are informative of changes between systems. Thus, Cdn can be viewed as implementing similar operations within an end-to-end, supervised framework.
2.4 Amortized causal inference
While the majority of causal discovery algorithms must be fit or run from scratch for each data distribution, amortized causal discovery algorithms have addressed this task as a supervised machine learning problem. Using large numbers of synthetic datasets, generated by synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Li et al., 2020; Ke et al., 2022; Lorch et al., 2022; Petersen et al., 2023; Wu et al., 2024). Compared to traditional approaches, these algorithms trade explicit assumptions for implicit characteristics of the data simulation process (Montagna et al., 2024). These models are also fast and robust to noise. In this work, we simulate interventions that reflect biological perturbations, and our model architecture builds upon that of Wu et al. (2024), as it scales easily to hundreds of variables.
@@ 3 Methods
We introduce causal differential networks (Cdn), a causality inspired method for predicting intervention targets. On a high level, Cdn is motivated by the idea that differences in observations $D_{\text{obs}}$ and $D_{\text{int}}$ can be explained by the changes to their underlying causal mechanisms (Figure 1B).
Our model is composed of two modules: a causal structure learner (Section 3.1), and a differential network classifier (Section 3.2). Given a single dataset $D$, the causal structure learner predicts a putative $G$. The pair of datasets $D_{\text{obs}}$ and $D_{\text{int}}$ yields graphs $G_{\text{obs}}$ and $G_{\text{int}}$, whose features are input to the differential network, to predict targets $I$ (Figure 1C). Both modules are trained jointly, supervised by classification losses for the graph and predicted targets. Finally, individual samples are known to be noisy in single-cell biology and are often analyzed in “pseudo-bulk” (Blake et al., 2003). Thus, instead of operating over raw data, we also represent datasets in terms of their summary statistics.
Figure 2: Our model is composed of a causal structure learner and a differential network. Learned components in yellow (causal structure learner) and green (differential network). Both modules are trained jointly to optimize $\mathcal{L}=\mathcal{L}_{G}+\mathcal{L}_{I}$.
3.1 Causal structure learner
The causal structure learner takes as input dataset $D$ and outputs the predicted $G$. In principle, this module can be implemented with any data-efficient causal discovery model. Here, we based the architecture on Sea (Wu et al., 2024), as its runtime complexity does not explicitly depend on the number of samples (which can be large for $D_{\text{obs}}$).
Input featurization  We represent dataset $D$ through two types of statistics. First, we compute global statistics $\rho\in\mathbb{R}^{N\times N}$, like pairwise correlation, over all variables. Second, we obtain estimates of local causal structure by running classical causal discovery algorithms like FCI (Spirtes et al., 1995) over small subsets of variables. We denote the set of estimates as $E^{\prime}\in\mathcal{E}^{k\times k\times T}$, where $\mathcal{E}$ is the space of possible edge types (e.g. ,$\leftrightarrow$,-$\rightarrow$ , - , →), and $k,T$ are hyperparameters that trade off accuracy vs. runtime.
To compare $\rho$ and $E^{\prime}$ in the same space, we linearly project elements of $\rho$ from $\mathbb{R}$ to $\mathbb{R}^{d}$, and map categorical edge types into $\mathbb{R}^{d}$ using a learned embedding, i.e. like token embeddings in language models (Devlin et al., 2019). Thus, the dataset $D$ is summarized into a matrix of size $\mathbb{R}^{N\times N\times T\times d}$.
Model architecture  Both the causal structure learner and differential network are implemented with a series of attention-based blocks. Since scaled dot-product attention scales quadratically as sequence length (Vaswani et al., 2017), we attend over one dimension at a time. For example, given a $N^{2}$ entries in an adjacency matrix, naive self-attention would cost $O(N^{4})$ operations. Instead, we apply self-attention along all nodes in the “outgoing edge” direction, with the “incoming edges” as a batch dimension, followed by the opposite – resulting in two operations, each of cost $O(N\times N^{2})$. Following Rao et al. (2021), we use pre-layer normalization on each self-attention, followed by dropout and residual connections:
$h\leftarrow h+\text{Dropout}(\text{Self-Attn}(\text{LayerNorm}(h)))$ (4)
where $h\in\mathbb{R}^{\dots\times d}$ denotes an intermediate hidden representation. This design takes after Transformer models for higher dimensional data in other domains, e.g. Axial Transformers for videos (Ho et al., 2020) and MSA Transformers for aligned protein sequences (Rao et al., 2021). A key difference is the graph modality: when making edge or node-level predictions, the model should be equivariant to node labeling. This is accomplished by adding randomly-permuted “positional” embeddings to each node and edge representation. For more details, please see Appendix A.
Edge-level outputs  Let $h_{\text{obs}},h_{\text{int}}\in\mathbb{R}^{N\times N\times d}$ denote the final layer representations of each dataset, after pooling over the $T$ dimension. These representations are input to the differential network. In addition, we introduce a graph prediction head,
$z_{i,j}$ $=\text{FFN}([h_{i,j},h_{j,i}])$ (logits) (5)
$P(E_{i,j}=e)$ $=\sigma(z_{i,j})_{e}$ (softmax) (6)
where $\sigma$ denotes softmax over the set of output edge types (forwards, backwards, or no edge). The graph prediction loss $\mathcal{L}_{G}$ is simply the cross entropy loss between the predicted and true edge types, summed over the entire graph.
3.2 Differential network
Given a pair of graph representations $h_{\text{obs}},h_{\text{int}}$, the differential network predicts $I$, the set of intervention targets.
Input featurization  In addition to pairwise representations $h$, we compute node-level statistics (mean, variance), which are mapped to $\mathbb{R}^{d}$ via a feed-forward network, and concatenated along “incoming edge” dimension. This results in graph representation $h^{\prime}\in\mathbb{R}^{N\times(N+1)\times d}$.
We combine the graph representations either through concatenation or subtraction:
$h_{\text{cat}}$ $=[h^{\prime}_{\text{obs}},h^{\prime}_{\text{int}}]\in\mathbb{R}^{N\times(N+1)\times 2d}$ (7)
$h_{\text{diff}}$ $=h^{\prime}_{\text{int}}-h^{\prime}_{\text{obs}}\in\mathbb{R}^{N\times(N+1)\times d}.$ (8)
The former can express richer relationships between the two graphs, while the latter is more efficient on large graphs.
Model architecture  The differential network closely follows the architecture of the causal structure learner. However, since paired graph representation lacks the $T$ dimension (introduced by marginal estimates $E^{\prime}$), the model only needs to attend over two “length” dimensions (rows and columns of adjacency matrix). After the attention layers, we mean over all “incoming edges” and linearly project to binary node-level predictions, where a normalized 1 indicates an intervention target, and 0 indicates otherwise.
The target prediction loss $\mathcal{L}_{I}$ is the binary cross entropy loss between the predicted and true targets, computed independently for each node. This is to accommodate interventions with multiple targets, e.g. drugs with off-target effects.
Remark  We take an amortized causal discovery approach because it is hard to specify what assumptions a biological system ought to follow. This makes it difficult to provide identifiability guarantees without simplifying aspects of either the data or modeling framework. Instead, in Appendix C, we show that the differential network is a well-specified class of models. That is, there exist parametrizations of its attention architecture that map predicted graphs and global statistics (for hard and soft interventions) to the set of targets.
3.3 Implementation details
Training procedure  The overall training loss is $\mathcal{L}=\mathcal{L}_{G}+\mathcal{L}_{I}$ with L2 weight regularization. Both $\mathcal{L}_{G}$ and $\mathcal{L}_{I}$ supervise the causal structure learner, while only the latter supervises the differential network. We use the synthetic and real datasets as follows.